# A General Bayesian Algorithm for the Autonomous Alignment of Beamlines

### v1

@@ I Introduction
Some attempts at beamline alignment apply methods like genetic and differential evolution [6; 7; 8; 9], attempt to match beamline data to an online model [10; 11], or use families of commonly-used optimization algorithms [12]. These approaches are limited in that they give no guarantee of convergence to a global optimum. They also make no consideration of minimizing the number of function evaluations, and beamline optimization almost universally involves a prohibitively expensive-to-sample function, both at the real beamline (relying on the movement of precise motors, which can be slow) and on simulated digital twins (relying on computationally intensive ray-tracing [13] or Fourier-based methods [14]) meaning that their use is intractable for large numbers of dimensions. In contrast to the classical methods above, algorithms based on machine learning construct and fit a model to understand the effects of changing the parameter inputs, as well as the interaction of the output beam qualities (e.g. flux, spatial resolution, energy resolution, polarization, coherence), leading to a more efficient search of the parameter space. Some machine learning methods like reinforcement learning [15] suffer from similar drawbacks to the methods above in that they may take too long to learn enough to be useful, at which point beamline parameters and hyperparameters may have drifted substantially. From a practical point of view, then, we should greatly prefer alignment methods that converge as quickly as possible, and rely on little to no prior input.
@@
An ML framework well-suited for expensive-to-sample functions is Bayesian optimization, which performs well with no prior information on optimization problems that are expensive-to-sample, high-dimensional, and potentially very noisy. Bayesian optimization has been applied in such a wide variety of contexts as synchrotron light sources [16; 17], free-electron lasers [18], particle colliders [19], and laser–plasma-based ion sources [20]. These implementations, however, are typically applicable to single experiments; indeed, much of the difficulty in implementing machine learning solutions to any problem is the trade-off of specificity and generality where an algorithm that is specific enough to be effective in some context is too specific to be applied generally.
@@
Bayesian optimization is highly generalizable in the choice of the kernel model used to describe the parameter space, and that many facilities are moving toward shared software environments and shared data acquisition protocols like Bluesky [21; 22] suggest the benefit of a general agent. This paper demonstrates an implementation of a Bayesian agent that can learn the dynamics and idiosyncrasies of a particular beamline and can thus be deployed across many different beamlines with relatively little implementation cost by applying the same code to a range of optimization problems at different synchrotron and non-synchrotron facilities.
@@
In Sections II, III, and IV we present a general but brief formulation of multiobjective Bayesian optimization with Gaussian process regression as it relates to this work (for a more thorough introduction see Frazier 23). Section V addresses beamline-specific considerations for Bayesian optimization, and Section VI presents their implementation in a software package. Section VII describes the application of the code at beamlines across the National Synchrotron Light Source II and Accelerator Test Facility at Brookhaven National Laboratory and the Advanced Light Source at Lawrence Berkeley National Laboratory, and presents the results of benchmarking on a simulated digital twin. Finally, Section VIII discusses the future development of the algorithm.
@@ IV.3 Multiobjective optimization
Optimization problems often require managing trade-offs. For example, a common beamline design consists of a secondary-source aperture (SSA), which cuts off some flux in the interest of having a smaller and tighter beam. One method of multiobjective optimization is scalarization, which is to use a function that maps a vector output to a scalar output, leading to one quantity to be maximized. In this work, we use affine scalarizations (i.e. assigning a weight to each objective and summing them) to construct a single fitness function over which to optimize an acquisition function, though there are other useful ways to carry out multiobjective optimization like Pareto efficient searches (i.e. finding the set of inputs where no one objective can be increased without decreasing some other objective) or thresholding (i.e. finding the set of points where all objectives meet some criteria).
@@ VI.1 Beamline Optimization (Blop)
Our beamline alignment tools are implemented in the Blop [^10] See https://nsls-ii.github.io/blop for documentation and tutorials. Python package, relying on the BoTorch Python package [30]. In Blop we develop a customized kernel which fits to latent beamline dimensions outlined in Section V.1, and weight common acquisition functions by the probabilistic constraint outlined in Section V.2. We also use BoTorch for model fitting and acquisition function optimization. The algorithm is used in terms of an agent, which we instantiate with motors and diagnostic equipment. We can “tell” the agent about the values of pre-defined objectives (e.g. beam height, coherence) and “ask” it for new points to sample. The agent wraps the steps of Bayesian optimization in a .learn() method, which yields a plan accepted by the Bluesky experiment orchestration system.
@@ VI.2 Bluesky
Bluesky [21; 22] is a software package that allows for the orchestration and execution of experiments from Python, and is in the process of being adapted by various light sources. [^11] See https://blueskyproject.io/ We design Blop with Bluesky in mind, as it can use Bluesky to automatically take data, analyze it, and optimize the inputs with the same feedback and control systems used for beamline experiments. This allows the Blop agent to both command and control the beamline, leading to an easier implementation. [^12] Though Blop is not limited to Bluesky facilities and can be made to only “command” the experiments using only its “ask” and “tell” methods. Bluesky has been mainly developed by NSLS-II, with a growing international collaboration at multiple facilities where it is used and expanded. The adoption of a single standard for experimental control and analysis across many facilities allows us to apply the same automated alignment tools with relatively little effort.
@@ VII.2 Alignment of a Johann Spectrometer at the ISS beamline
The manual optimization is rendered difficult by the interplay between the toroidal mirror angle, monochromator height and angle, all of them changing the beam height and interfering with 4-jaw slits. With the use of the described automated alignment, we were able to maximize the power density on the sample in under 5 minutes, with a final beam size of 1 mm x 0.3 mm (HxV, FWHM), close to the theoretical limit calculated by raytracing (Fig. 8) with an improvement of the intensity by over factor 2 over our best effort using manual alignment.
Figure 8: Four different beam configurations at the Advanced Light Source’s Beamline 5.3.1 during automated alignment. In total, the photon transport has 4 active degrees of freedom: the focusing mirror pitch and tangential bend, the channel cut crystal angle and height. The upper left panel shows the initial, manually aligned beam and the lower right the final beam after automated alignment. Upper right and lower left panels show intermediary points collected in the automated alignment process.
@@ VII.4 Alignment of an electron beam at the Accelerator Test Facility
The Accelerator Test Facility (ATF) is a user facility at Brookhaven National Laboratory offering the combination of an 80 MeV electron beam synchronized with a Terawatt picosecond carbon dioxide laser [34]. This gives it the capability to develop cutting-edge electron beam techniques including ultrafast electron diffraction and microscopy [35], as well as free-electron laser techniques including direct laser acceleration and using Compton scattering as a high energy x-ray source [36]. We modulate three bending quadrupole electromagnets and a solenoid to manipulate the shape of the beam, for a total of 4 degrees of freedom.
@@ VII.5 Simulated alignment of the TES beamline
Figure 10: The 8-dimensional optimization of the simulated TES beamline, where the degrees of freedom comprise the toroidal and Kirkpatrick-Baez mirrors. The colors show different varieties of Bayesian optimization algorithms with and without both of latent inputs and composite outputs, with both the cumulative maximum of all individual runs (thin lines) and the median cumulative maximum (thick line). Each variety starts out with a quasi-random sampling of 32 points, and then performs a Bayesian optimization loop with the expected improvement acquisition function. The benefit of using both latent inputs and composite outputs is shown, as we can achieve a better optimum more robustly and more quickly.
Figure 11: The 4-dimensional optimization of just the K-B mirrors, whose motors are each misaligned by up 0.05 mm. After an initial quasi-random sample of 16 points, the agent is able to almost instantly return to the optimal alignment.
@@
For benchmarking, we consider the digital twin of the TES beamline at NSLS-II. In the 8-dimensional case, we use the six degrees of freedom outlined in Section VII.1, but also allow the toroidal mirror to yaw and translate horizontally for a total of 8 degrees of freedom. Each K-B motor can move up to $\pm$ 0.25 mm from a fiducial starting point, while the range of each toroidal motor was bounded by the points where the misalignment of that motor caused the flux through the SSA to fall to 50% of the maximum. The results of this benchmark are shown in Figure 10.

### v2

@@ I Introduction
Some attempts at beamline alignment apply methods like genetic and differential evolution [6; 7; 8; 9], attempt to match beamline data to an online model [10; 11], or use families of commonly-used optimization algorithms [12; 13]. These approaches are limited in that they give no guarantee of convergence to a global optimum. They also make no consideration of minimizing the number of function evaluations, and beamline optimization almost universally involves a prohibitively expensive-to-sample function, both at the real beamline (relying on the movement of precise motors, which can be slow) and on simulated digital twins (relying on computationally intensive ray-tracing [14] or Fourier-based methods [15]) meaning that their use is intractable for large numbers of dimensions. In contrast to the classical methods above, algorithms based on machine learning construct and fit a model to understand the effects of changing the parameter inputs, as well as the interaction of the output beam qualities (e.g. flux, spatial resolution, energy resolution, polarization, coherence), leading to a more efficient search of the parameter space. Some machine learning methods like reinforcement learning [16] suffer from similar drawbacks to the methods above in that they may take too long to learn enough to be useful, at which point beamline parameters and hyperparameters may have drifted substantially. From a practical point of view, then, we should greatly prefer alignment methods that converge as quickly as possible, and rely on little to no prior input.
@@
A machine learning framework well-suited for expensive-to-sample functions is Bayesian optimization, which performs well with no prior information on optimization problems that are expensive-to-sample, high-dimensional, and potentially very noisy. Bayesian optimization has been applied in such a wide variety of contexts as synchrotron light sources [17; 18], free-electron lasers [19], particle colliders [20], and laser–plasma-based ion sources [21]. These implementations, however, are typically applicable to single experiments; indeed, much of the difficulty in implementing machine learning solutions to any problem is the trade-off of specificity and generality where an algorithm that is specific enough to be effective in some context is too specific to be applied generally.
@@
Bayesian optimization is highly generalizable in the choice of the kernel model used to describe the parameter space, and that many facilities are moving toward shared software environments and shared data acquisition protocols like Bluesky [22; 23] suggest the benefit of a general agent. This paper demonstrates an implementation of a Bayesian agent that can learn the dynamics and idiosyncrasies of a particular beamline and can thus be deployed across many different beamlines with relatively little implementation cost by applying the same code to a range of optimization problems at different synchrotron and non-synchrotron facilities.
@@
In Sections II, III, and IV we present a general but brief formulation of multiobjective Bayesian optimization with Gaussian process regression as it relates to this work (for a more thorough introduction see Frazier 24). Section V addresses beamline-specific considerations for Bayesian optimization, and Section VI presents their implementation in a software package. Section VII describes the application of the code at beamlines across the National Synchrotron Light Source II and Accelerator Test Facility at Brookhaven National Laboratory and the Advanced Light Source at Lawrence Berkeley National Laboratory, and presents the results of benchmarking on a simulated digital twin. Finally, Section VIII discusses the future development of the algorithm.
@@ IV.3 Multiobjective optimization
Optimization problems often require managing trade-offs. For example, a common beamline design consists of a secondary-source aperture (SSA), which cuts off some flux in the interest of having a smaller and tighter beam. One method of multiobjective optimization is scalarization, which is to use a function that maps a vector output to a scalar output, leading to one quantity to be maximized. In this work, we use affine scalarizations (i.e. assigning a weight to each objective and summing them) to construct a single fitness function over which to optimize an acquisition function.
There are, however, are other useful ways to carry out multiobjective optimization like Pareto efficient searches (i.e. finding the set of inputs where no one objective can be increased without decreasing some other objective). But while fully multiobjective methods do allow for more flexibility in the alignment, that flexibility is not compatible with an autonomous beamline, which must decide on a single best beam and thus must collapse the beam to a single fitness function.
@@ VI.1 Beamline Optimization (Blop)
Our beamline alignment tools are implemented in the Blop [^10] See https://nsls-ii.github.io/blop for documentation and tutorials. Python package, relying on the BoTorch Python package [31]. In Blop we develop a customized kernel which fits to latent beamline dimensions outlined in Section V.1, and weight common acquisition functions by the probabilistic constraint outlined in Section V.2. We also use BoTorch for model fitting and acquisition function optimization. The algorithm is used in terms of an agent, which we instantiate with motors and diagnostic equipment. We can “tell” the agent about the values of pre-defined objectives (e.g. beam height, coherence) and “ask” it for new points to sample. The agent wraps the steps of Bayesian optimization in a single customizable routine (implemented in Python as a .learn() method), which yields a plan accepted by the Bluesky experiment orchestration system. [^11] In effect, a single button that can be pressed to align the beamline. This routine can be tailored to each beamline (or to each alignment problem for a given beamline). Encapsulating the optimization as a single process simplifies the alignment from the point of view of the user, making experimentation more accessible to users who have less familiarity with a given beamline (i.e. hardware, data acquisition, control systems), or with less familiarity with software in general.
@@ VI.2 Bluesky
Bluesky [22; 23] is a software package that allows for the orchestration and execution of experiments from Python, and is in the process of being adapted by various light sources. [^12] See https://blueskyproject.io/ We design Blop with Bluesky in mind, as it can use Bluesky to automatically take data, analyze it, and optimize the inputs with the same feedback and control systems used for beamline experiments. This allows the Blop agent to both command and control the beamline, leading to an easier implementation. [^13] Though Blop is not limited to Bluesky facilities and can be made to only “command” the experiments using only its “ask” and “tell” methods. Bluesky has been mainly developed by NSLS-II, with a growing international collaboration at multiple facilities where it is used and expanded. The adoption of a single standard for experimental control and analysis across many facilities allows us to apply the same automated alignment tools with relatively little effort.
@@ VII.2 Alignment of a Johann Spectrometer at the ISS beamline
The manual optimization is rendered difficult by the interplay between the toroidal mirror angle, monochromator height and angle, all of them changing the beam height and interfering with 4-jaw slits. With the use of the described automated alignment, we were able to maximize the power density on the sample in under five minutes, with a final beam size of 1 mm x 0.3 mm (HxV, FWHM), close to the theoretical limit calculated by raytracing (Figure 8) with an improvement of the intensity by over a factor of two over our best effort using manual alignment.
Figure 8: Four different beam configurations at the Advanced Light Source’s Beamline 5.3.1 during automated alignment. In total, the photon transport has four active degrees of freedom: the focusing mirror pitch and tangential bend, the channel cut crystal angle and height. The upper left panel shows the initial, manually aligned beam and the lower right the final beam after automated alignment. Upper right and lower left panels show intermediary points collected in the automated alignment process.
@@ VII.4 Alignment of an electron beam at the Accelerator Test Facility
The Accelerator Test Facility (ATF) is a user facility at Brookhaven National Laboratory offering the combination of an 80 MeV electron beam synchronized with a Terawatt picosecond carbon dioxide laser [35]. This gives it the capability to develop cutting-edge electron beam techniques including ultrafast electron diffraction and microscopy [36], as well as free-electron laser techniques including direct laser acceleration and using Compton scattering as a high energy x-ray source [37]. We modulate three bending quadrupole electromagnets and a solenoid to manipulate the shape of the beam, for a total of four degrees of freedom.
@@ VII.5 Simulated alignment of the TES beamline
Figure 10: The 8-dimensional optimization of the simulated TES beamline, where the degrees of freedom comprise the toroidal and Kirkpatrick-Baez mirrors. The colors show different varieties of Bayesian optimization algorithms with and without both of latent inputs and composite outputs, with both the cumulative maximum of all individual runs (thin lines) and the median cumulative maximum (thick line). Each variety starts out with a quasi-random sampling of 32 points (shaded light blue), and then performs a Bayesian optimization loop with the expected improvement acquisition function. The benefit of using both latent inputs and composite outputs is shown, as we can achieve a better optimum more robustly and more quickly.
Figure 11: The 4-dimensional optimization of just the K-B mirrors, whose motors are each misaligned by up 0.05 mm. After an initial quasi-random sample of 16 points (shaded light blue), the agent is able to almost instantly return to the optimal alignment.
@@
For benchmarking, we consider the digital twin of the TES beamline at NSLS-II. In the 8-dimensional case, we use the six degrees of freedom outlined in Section VII.1, but also allow the toroidal mirror to yaw and translate horizontally for a total of eight degrees of freedom. Each K-B motor can move up to $\pm$ 0.25 mm from a fiducial starting point, while the range of each toroidal motor was bounded by the points where the misalignment of that motor caused the flux through the SSA to fall to 50% of the maximum. The results of this benchmark are shown in Figure 10.